OpenStudy (anonymous):

(-2/3)^-2 (simplify)

OpenStudy (anonymous):

9/4

jimthompson5910 (jim_thompson5910):

\[\Large \left(\frac{2}{3}\right)^{-2}\] \[\Large \left(\frac{3}{2}\right)^{2}\] \[\Large \left(\frac{-3}{2}\right)^{2}\] \[\Large \frac{(-3)^2}{2^2}\] \[\Large \frac{9}{4}\] So \[\Large \left(\frac{2}{3}\right)^{-2} = \frac{9}{4}\]